OpenStudy (anonymous):

Please help :) Find the points of intersection of the graphs of y = x^2 and y=3x-2

OpenStudy (campbell_st):

equate the 2 equations and solve \[x^2 = 3x -2\] or \[x^2 - 3x + 2 = 0\] you'll need 2 points of intesection. So solve for x then substitute them into either equation to find the corresponding y values

OpenStudy (polaris_s0i):

points of intersection are when both graphs go through the same point. This means both X and Y in both functions are equal to each other. since Y of the first function and Y of the second function are equal, then set both equations equal to each other and solve it. x^2 = 3x -2 then move it over x^2 -3x + 2 = 0 and factor (x-2)(x-1) = 0 or x = 2 and x = 1

OpenStudy (anonymous):

thanks guys :) @campbell_st @polaris_s0i